# Transfer learning of state-based potential games for process optimization in decentralized manufacturing systems

### v1

@@ 6.3 Results on the BGS
During our experiments, we discovered that the TL-SbPG methods were able to achieve comparable performance to the native SbPG within a shorter time frame. Additionally, we observed that as the training continued uninterrupted for the full 9 training episodes, the performance of the transfer learning methods improved even further. We claim that the players could explore more state-action combinations. However, we note that the standard SbPG approach provided an optimal policy with respect to avoiding overflow and meeting production demands, and thus the power consumption was the primary focus for achieving convergence with reduced consumption.
Fig. 8 illustrates the power consumption and potential value during the training and testing processes. The graph demonstrates a consistent and significant reduction in power consumption over time across each method. This reduction in power consumption has a direct impact on the global multi-objective of the production system, as outlined in Eq. (25). By achieving a notable decrease in power consumption, the overall performance and efficiency of the production system as well as the potential value are improved, aligning with the system’s objectives.
(a) Power consumption
(b) Potential value
@@
(a) SW approach
(b) MOM approach
@@
In Fig. 10, we present a comparison of the overall performances between each TL-SbPG approach and the baseline. The results show that the TL-SbPG approaches outperform the baseline in two aspects. First, the power consumption is significantly reduced ranging up to 12.46%. Second, the potential value is enhanced by approximately between 22.3% and 47.2% compared to the baseline.
@@
Additionally, we tune and identify the optimal hyperparameter configuration using an automated multi-resolution grid search [38] to guarantee the optimal configuration of the learning algorithms. One of the hyperparameters is threshold $\beta_{\text{TF}}$, which plays a crucial role as a kick-off indicator of the transfer learning process, as explained in Eq. (20). For the TL-SbPG with pre-defined similarities, the best results are obtained when the threshold $\beta_{\text{TF}}$ for the MOM and SW approaches are 0.80 and 0.40, respectively. Meanwhile, for the TL-SbPG without pre-defined similarities, the threshold $\beta_{\text{TF}}$ for RBF networks is 0.24. Table II shows an ablation study of threshold $\beta_{\text{TF}}$ on the MOM and SW approaches.
@@ 6.4 Results on the LS-BGS
Figure 11: Similarity matrix of actuators in the LS-BGS.
@@ 7 Conclusions
In future research, we would like to develop other GT structures than SbPGs on larger-scale industrial control domains, e.g. Stackelberg games. Then, we will test the proposed transfer learning methods across different game structures. Additionally, we aim to enhance the transfer learning methods by developing a simpler approach that requires less communication between players to avoid noise and accelerate the transfer learning processes.
@@
This research did not receive specific funding from any grant provided by public, commercial, or not-for-profit sectors.

### v3

@@ 6.3 Results on the BGS
Through our experiments, we found that the TL-SbPG methods achieved comparable performance to the vanilla SbPG, but within a significantly shorter time frame, which requires approximately 2-4 fewer training episodes (or 22.2-44.4% less) to converge. Additionally, we observed that as the training continued uninterrupted for the full 9 training episodes, the performance of the transfer learning methods improved even further. We claim that the players could explore more state-action combinations. However, we note that the standard SbPG approach provided an optimal policy with respect to avoiding overflow and meeting production demands, and thus, the power consumption was the primary focus for achieving convergence with reduced consumption.
Fig. 8 illustrates the power consumption and potential value during the training and testing processes. The graph demonstrates a consistent and significant reduction in power consumption over time across each method. This reduction in power consumption has a direct impact on the global multi-objective of the production system, as outlined in Eq. (25). By achieving a notable decrease in power consumption, the overall performance and efficiency of the production system, as well as the potential value are improved, aligning with the system’s objectives.
(a) Power consumption over time (x-axis: simulation time in 100s, y-axis: power in kW).
(b) Potential value over time (x-axis: simulation time in 100s, y-axis: potential value).
@@
(a) SW approach, x-axis: fill-level of previous buffer (%), y-axis: fill-level of subsequent buffer (%).
(b) MOM approach, x-axis: fill-level of previous buffer (%), y-axis: fill-level of subsequent buffer (%).
@@
In Fig. 10, we present a comparison of the overall performances between each TL-SbPG approach and the baseline. The results show that the TL-SbPG approaches outperform the baseline in two aspects. First, the power consumption is significantly reduced, ranging up to 12.46%. Second, the potential value is enhanced by approximately between 22.3% and 47.2% compared to the baseline.
@@
Additionally, we tune and identify the optimal hyperparameter configuration using an automated multi-resolution grid search [48] to guarantee the optimal configuration of the learning algorithms. One of the hyperparameters is the threshold $\beta_{\text{TF}}$, which plays a crucial role as a kick-off indicator of the transfer learning process, as explained in Eq. (20). For the TL-SbPG with pre-defined similarities, the best results are obtained when the threshold $\beta_{\text{TF}}$ for the MOM and SW approaches are 0.80 and 0.40, respectively. Meanwhile, for the TL-SbPG without pre-defined similarities, the threshold $\beta_{\text{TF}}$ for RBF networks is 0.24. Table II shows an ablation study of threshold $\beta_{\text{TF}}$ on the MOM and SW approaches.
@@
Then, we validate the effectiveness of TL-SbPGs in arbitrary BGS environments, including serial-parallel processes and module additions. We follow the three distinct scenarios outlined in [21]. Instead of training the policies from scratch, we reuse the pre-trained policies from the default BGS setup and apply an additional 1-2 training episodes, depending on the scenario’s complexity. This approach, as demonstrated in [3], proves effective and highlights the system’s plug-and-play functionality. In serial-parallel processes with the additional module configuration of 1-3-2//3-4, the transfer learning process shifts from involving both vacuum pumps to focusing on the actuators within the duplicated Module 3 because they are identical. Table III provides a comparative summary of the TL-SbPG results across different scenarios, which shows that the proposed approaches consistently outperform the vanilla SbPG baseline, with MOM and RBF delivering the best performance compared to the SW approach.
Table III: Comparison between SbPGs and TL-SbPGs on the arbitrary environments of the BGS.
Algorithm
Demand
(L/s)
Overflow
(L/s)
Power
(kW/s)
Potential
Sequence: 1-3-2-4
Baseline 0.000 0.000 0.479 11.960
SW 0.000 0.000 0.465 12.065
MOM 0.000 0.000 0.460 12.897
RBF 0.000 0.000 0.460 12.844
Sequence: 1-2//3-4
Baseline 0.000 0.000 0.488 11.883
SW 0.000 0.000 0.471 11.997
MOM 0.000 0.000 0.465 12.765
RBF 0.000 0.000 0.467 12.800
Sequence: 1-3-2//3-4
Baseline 0.000 0.012 0.669 15.410
SW 0.000 0.008 0.615 15.984
MOM 0.000 0.001 0.603 17.010
RBF 0.000 0.002 0.612 16.867
We conducted a paired t-test to compare the utility values of TL-SbPGs and SbPGs across 12 experiments on the BGS under varied setups and TL-SbPGs variances. The results showed that TL-SbPGs significantly outperformed SbPGs, with a mean utility improvement of 1.51 ($t(11)=3.70,p=0.0035$). This confirms that TL-SbPGs provide consistently better performance under diverse conditions.
Finally, we modified the power consumption and control parameters of the vacuum pumps in the default setting of the BGS, representing Player 2 and Player 4, both involved in the transfer learning process. Specifically, we reduced Player 2’s transport capability by 25% and increased its power consumption by 15%, while for Player 4, we increased the transport capability by 10% and raised the power consumption by 25%. Due to these changes in control parameters, unlike the previous ablation study, we retrained the system from scratch. The results, shown in Table IV, demonstrate that the proposed approach again significantly outperforms the baseline. The findings from the last two experiments confirm that TL-SbPG remains effective across varying environments and control parameters, further supporting the validity of our convergence proof and theorem in this study.
Table IV: Comparison between SbPGs and TL-SbPGs on the BGS’ performances with modified control and power consumption parameters of Player 2 and 4.
Algorithm
Demand
(L/s)
Overflow
(L/s)
Power
(kW/s)
Potential
Baseline 0.000 0.000 0.710 6.741
SW 0.000 0.000 0.629 8.784
MOM 0.000 0.000 0.618 8.901
RBF 0.000 0.000 0.621 8.775
@@ 6.4 Results on the LS-BGS
Figure 11: Similarity matrix of actuators in the LS-BGS. The x- and y-axes represent player indices, and the color scale indicates the degree of similarity between each pair of players.
@@
Furthermore, we demonstrate that the number of players involved in transfer learning is not naturally limited. To this end, we conduct simulations with four different configurations of player involvement. Table V presents the results for these varying player numbers using the MOM approach, which indicates that six players are the most effective in the LS-BGS. In the arbitrary environment experiments of the BGS, it is noted that involving four players in the 1-3-2//3-4 sequence is more advantageous than involving only two players. This suggests that the number of players is not restricted and can be increased as long as there are sufficient similar players available.
Table V: Ablation study of the number of players involved in the transfer learning process in the LS-BGS.
Algorithm
Demand
(L/s)
Overflow
(L/s)
Power
(kW/s)
Potential
Baseline 0.000 0.011 1.183 37.951
2P : P5-P10
MOM 0.000 0.098 1.087 42.703
4P : P5-P10, P2-P6
MOM 0.000 0.071 1.052 44.940
6P: P5-P10, P2-P6, P1-P4
MOM 0.000 0.065 1.049 45.634
8P: P5-P10, P2-P6, P1-P4, P3-P8
MOM -0.001 0.101 1.060 41.330
From this ablation study, we observe that the distributed learning architecture imposes no limitation on the number of players in either the vanilla SbPG or TL-SbPG frameworks. The number of players involved in the transfer learning process is determined solely by the availability and similarity of players within the system. To be noted, the number of inputs to the learning algorithm is dependent on the number of neighbouring players. Consequently, the complexity of the system scales only with the number of neighbouring players, which is typically quite limited in production processes.
We then extend our experimental setup of the LS-BGS to include serial-parallel processes, as illustrated in Fig. 13. Similar to the plug-and-play functionalities in the BGS, we can reuse the trained policies from the default sequential setup and retrain them with an additional three training episodes. We also maintain the same number of players involved in the transfer learning process, as determined from the previous ablation study. The results, presented in Table VI, indicate that the LS-BGS consistently outperforms the vanilla SbPG, while the RBF model does not significantly enhance the system’s performance compared to the SW and MOM approaches. These findings demonstrate that TL-SbPGs exhibit robustness and scalability in more complex and varied environments, further validating our proof of convergence.
Figure 13: Schematic of the LS-BGS with serial-parallel processes.
Table VI: Comparison between SbPGs and TL-SbPGs on the LS-BGS’ performances with serial-parallel processes.
Algorithm
Demand
(L/s)
Overflow
(L/s)
Power
(kW/s)
Potential
Baseline -0.001 0.169 1.479 35.406
SW 0.000 0.070 1.285 38.957
MOM 0.000 0.066 1.243 39.035
RBF -0.009 0.187 1.370 35.980
A paired t-test was conducted to evaluate the performance of TL-SbPGs against SbPGs across six experiments in the LS-BGS, involving both sequential and serial-parallel processes with varied TL-SbPG variances. The analysis revealed a statistically significant performance advantage for TL-SbPG, with a mean utility improvement of 2.27 ($t(5)=3.77,p=0.0133$). These results indicate once again that TL-SbPGs offers superiority across different LS-BGS process structures.
After discussing the advantages of our proposed approaches in two experimental setups, we outline the limitations identified through our findings. The experiments revealed three limitations, although these do not diminish the overall contributions of the proposed TL-SbPGs approach. First, the method requires a digital representation of the system to train the players’ policies due to the randomness during exploration, which makes it unsafe to train directly on real systems. This limitation can be addressed using a model-based GT approach, as suggested in [21, 44]. Second, the SbPG utilizes best-response learning, which relies on ad-hoc random sampling for exploration. During periods of high exploration, we disabled the transfer learning mechanism. If a smoother sampling method were utilized for exploration, transfer learning could potentially be applied from the start, which leads to faster learning. Third, training the systems requires a large number of system interactions, which can be costly. This issue can also be mitigated through the model-based GT.
@@ 7 Conclusions
In future research, we aim to implement the proposed transfer learning methods across various game structures, including distributed Stackelberg strategies in SbPGs [57]. Additionally, we plan to integrate TL-SbPGs into the model-based GT domain, which enables learning within virtual environments. This would reduce the number of system interactions required and eliminate the need for a digital representation of the system. Additionally, we aim to enhance the transfer learning methods by developing a simpler approach that requires less communication between players to avoid noise and accelerate the transfer learning processes.
@@
This research did not receive specific funding from any grant provided by the public, commercial, or not-for-profit sectors.